# A Layered Protocol Architecture for the Internet of Agents

### v2

@@ 6.2. Implementation Strategy
    Extend A2A with L9 Capabilities: A2A’s extension mechanism could incorporate SNL handshake messages (SNL-HELLO, SNL-SELECT, SNL-LOCK) as a standardized extension, enabling semantic negotiation while maintaining backward compatibility.
@@
    Prototype L9 SNL: Develop reference implementations of the Semantic Negotiation Layer as libraries that work with existing protocols, demonstrating semantic validation, context negotiation, and integration with LLM-based agents.
@@ 6.3. Remaining Challenges
  • Performance Overhead: The semantic handshake introduces latency. This must be amortized by using persistent ”semantic sessions” that survive across multiple L8 message exchanges.
  • The Bootstrapping Problem: How is the initial registry of Shared Contexts created? This requires cross-industry standardization efforts, similar to how schema.org emerged for web content. Early domain-specific pilots are essential.
  • Representation Alignment: If two agents use different embedding models, their vector spaces will not align. This remains an open research problem (sucholutsky2023, ). The SNL provides the protocol to negotiate this, but the underlying ML techniques must advance.
@@ 7. Conclusion
This paper proposes a two-layer extension building on existing work. The Agent Communication Layer (L8) formalizes structural capabilities emerging in A2A (a2a2024, ), MCP (mcp2024, ), and SLIM (slim2025, ), standardizing message envelopes, task delegation, and interaction patterns. The Agent Semantic Negotiation Layer (L9) introduces protocol-level semantic negotiation—enabling agents to discover, negotiate, and lock formal contexts defining concepts and parameters. This addresses a capability gap in current protocols, where semantic alignment occurs informally at the application layer.
Separating syntactic structure (L8) from semantic meaning (L9) supports both transactional patterns (agent orchestration) and coordinative patterns (agent collaboration). Challenges remain: semantic handshake latency, context registry standardization, semantic injection defense, and governance models for Schema Authorities. However, multi-agent systems are already deployed in supply chains, distributed systems, and collaborative reasoning. Building on existing protocols while adding formal semantic negotiation could provide improved interoperability and reliability for these emerging applications.

### v3

@@ 6.2. Implementation Strategy
    Extend A2A with L9 Capabilities: A2A’s extension mechanism could incorporate SL handshake messages (SL-HELLO, SL-SELECT, SL-LOCK) as a standardized extension, enabling semantic negotiation and **coordination** while maintaining backward compatibility.
@@
    Prototype L9 SL: Develop reference implementations of the Agent Semantic Layer as libraries that work with existing protocols, demonstrating semantic validation, context negotiation, and state consensus integration with LLM-based agents.
@@ 6.3. Remaining Challenges
  • Performance Overhead: The semantic handshake and consensus rounds introduce latency. This must be amortized by using persistent ”semantic sessions” that survive across multiple L8 message exchanges.
  • The Bootstrapping Problem: How is the initial registry of Shared Contexts created? This requires cross-industry standardization efforts, similar to how schema.org emerged for web content.
  • Representation Alignment: If two agents use different embedding models, their vector spaces will not align. This remains an open research problem (sucholutsky2023, ). The SL provides the protocol to negotiate this, but the underlying ML techniques must advance.
@@ 7. Conclusion
This paper proposes a two-layer extension building on existing work. The Agent Communication Layer (L8) formalizes structural capabilities emerging in A2A (a2a2024, ), MCP (mcp2024, ), and SLIM (slim2025, ), standardizing message envelopes, task delegation, and interaction patterns. The Agent Semantic Layer (L9) introduces protocol-level semantic negotiation and collective consensus—enabling agents to discover, negotiate, and lock formal contexts defining concepts and parameters, and to align system states across populations. This addresses a capability gap in current protocols, where semantic alignment and consensus occur informally at the application layer.
Separation of syntactic structure (L8) from semantic meaning and state (L9) supports both transactional patterns (agent orchestration) and coordination patterns (agent collaboration). Challenges remain: semantic handshake latency, context registry standardization, semantic injection defense, and governance models for Schema Authorities. However, multi-agent systems are already deployed in supply chains, distributed systems, and collaborative reasoning. Building on existing protocols while adding formal semantic coordination and consensus could provide improved interoperability and reliability for these emerging applications.